\begin{thebibliography}{40}

%\bibitem{Ara}D. Arapura,  {\it K$\ddot {\rm a}$hler solvmanifolds}, Int. Math. Res. Not. {\bf 3} (2004), 131--137, MR2038772, Zbl 1091.53047.
\bibitem{AN}D. Arapura, M. Nori,  {\it Solvable fundamental groups of algebraic varieties and K\"ahler manifolds}, Compositio Math. {\bf 116} (1999), no. 2, 173--188, MR1686777, Zbl 0971.14020.


%\bibitem{Aus}L. Auslander, {\it An exposition of the structure of solvmanifolds}, I. Algebraic theory.  Bull. Amer. Math. Soc. {\bf 79}  (1973), no. 2, 227--261, MR0486307, Zbl 0265.22016.

\bibitem{B}
O. Baues, {\it Infra-solvmanifolds and rigidity of subgroups in solvable linear algebraic groups}, Topology {\bf 43} (2004), no. 4, 903--924.%  MR2061212, Zbl 1059.57022.
%\bibitem{OB} O. Baues and B. Klopsch, Deformations and rigidity of lattices in solvable Lie groups. J. Topol. (online published)
%\bibitem{BD}Y. Benoist, K. Dekimpe, The uniqueness of polynomial crystallographic actions. Math. Ann. {\bf 322} (2002), no. 3, 563--571.
\bibitem{Bro} K. S. Brown, Cohomology of groups, Graduate Texts in Mathematics, vol. 87, Springer-Verlag, New York, 1982. 
%\bibitem{Brf}K. S. Brown, Homological criteria for finiteness. Comment. Math. Helv.{\bf  50} (1975), 129--135. 
%\bibitem{BC} O. Baues,  V. Cort\'es,{\it Aspherical K\"ahler manifolds with solvable fundamental group}, Geom. Dedicata {\bf 122} (2006), 215--229, MR2295551, Zbl 1128.53043.

%  \bibitem{BG}
%C. Benson, and C. S. Gordon, {\it K$\ddot {\rm a}$hler and symplectic structures on nilmanifolds}, 
%Topology {\bf 27} (1988), no. 4, 513--518, MR0976592, Zbl 0672.53036. 
% \bibitem{BGs} C. Benson, C. S. Gordon,{\it K\"ahler structures on compact solvmanifolds}, Proc. Amer. Math. Soc. {\bf 108} (1990), no. 4, 971--980. MR0993739, Zbl 0689.53036.
%\bibitem{Boc}Ch. Bock, On Rational Homotopy and Minimal Models.http://www.drchristophbock.de/ratHom.pdf
\bibitem{Bor}
A. Borel, Linear algebraic groups 2nd enl. ed Springer-Verlag (1991).
\bibitem{Cav} G. R. Cavalcanti,
The Lefschetz property, formality and blowing up in symplectic geometry. Trans. Amer. Math. Soc. {\bf 359} (2007), no. 1, 333--348.
\bibitem{CNY}
B. Cappelletti-Montano, A. De Nicola, and I. Yudin, Hard Lef- schetz Theorem for Sasakian manifolds, preprint, arxiv: 1306.2896, to appear in J. Differential Geom.
%\bibitem{BW}A. Borel, N. Wallach,  Continuous cohomology, discrete subgroups, and representations of reductive groups. Second edition. Mathematical Surveys and Monographs, {\bf 67}. American Mathematical Society, Providence, RI, 2000.
%\bibitem{De} D. Burde, K. Dekimpe, K. Vercammen, Complete LR-structures on solvable Lie algebras. J. Group Theory 13 (2010), no. 5, 703--719.
%\bibitem{Deki}K. Dekimpe, P. Igodt,  Polycyclic-by-finite groups admit a bounded-degree polynomial structure. Invent. Math. {\bf 129} (1997), no. 1, 121–140.
%\bibitem{Dek} K. Dekimpe,  Solvable Lie algebras, Lie groups and polynomial structures.  Compositio Math. {\bf 121} (2000), no. 2,  183-204.
\bibitem{DGMS}P. Deligne, P. Griffiths, J. Morgan, and D. Sullivan, Real homotopy theory of Kahler manifolds. Invent. Math. {\bf 29} (1975), no. 3, 245--274.
%\bibitem{DER} N. Dungey, A. F. M. ter Elst, D. W. Robinson,  Analysis on Lie Groups with Polynomial Growth, Birkh\"auser (2003).
%\bibitem{FEH}Y. F$\acute {\rm e}$lix, S.  Halperin,  J-C. Thomas,  Rational homotopy theory. Graduate Texts in Mathematics, {\bf 205}. Springer-Verlag, New York, 2001.
%\bibitem{Fe}Y. F$\acute {\rm e}$lix, J. Oprea and D. Tanr$\acute {\rm e}$, Algebraic Models in Geometry, Oxford Graduate Texts in Mathematics 17, Oxford University Press 2008.
\bibitem{GMo} P. Griffiths, J. Morgan,  Rational homotopy theory and differential forms. Second edition. Progress in Mathematics, {\bf 16}. Springer, New York, 2013. 
\bibitem{Gro}
M. Gromov, A topological technique for the construction of differential equations and in- equalities, Actes Congr. Internat. Math. (Nice 1970), vol. 2, 1971, pp. 221--225.

\bibitem{Hai} R. M. Hain, The Hodge de Rham theory of relative Malcev completion. Ann. Sci. Ecole Norm. Sup. (4) {\bf 31} (1998), no. 1, 47--92.
\bibitem{H} K. Hasegawa, Minimal models of nilmanifolds. Proc. Amer. Math. Soc. {\bf 106} (1989), no. 1, 65--71. 
%\bibitem{Hc} K. Hasegawa, A class of compact Kahlerian solvmanifolds and a general conjecture. Geom. Dedicata {\bf 78} (1999), no. 3, 253--258.
%\bibitem{Hn} K. Hasegawa, A note on compact solvmanifolds with K\"ahler structures. Osaka J. Math. {\bf 43} (2006), no. 1, 131--135.
\bibitem{Haia}R. Hain, Remarks on non-abelian cohomology of proalgebraic groups. J. Algebraic Geom. {\bf 22} (2013), no. 3, 581--598.
%\bibitem{Haima} R. Hain, M. Matsumoto, Weighted completion of Galois groups and Galois actions on the fundamental group of $P^{1}-\{0,1,\infty\}$. Compositio Math.  {\bf 139} (2003), no. 2, 119–167. 
\bibitem{Halp} S. Halperin,Lectures on minimal models. M\'em. Soc. Math. France (N.S.) No. 9-10 (1983). 
\bibitem{Hat}
A. Hatcher,  Algebraic topology. Cambridge University Press, Cambridge, 2002. 
%\bibitem{Hatt} A. Hattori, Spectral sequence in the de Rham cohomology of fibre bundles. J. Fac. Sci. Univ. Tokyo Sect. I {\bf 8} 1960 289--331 (1960). 

\bibitem{Hoc} G. Hochschild,  Cohomology of algebraic linear groups. Illinois J. Math. {\bf 5} 1961 492--519. 
\bibitem{Hoi}  G. Hochschild, 
Rationally injective modules for algebraic linear groups. 
Proc. Amer. Math. Soc. {\bf 14} 1963 880--883. 
%\bibitem{HM} G. Hochschild, G. D. Mostow,  Cohomology of Lie groups. Illinois J. Math. {\bf 6} 1962 367--401.
%\bibitem{HS}G. Hochschild,  J.-P. Serre, Cohomology of group extensions. Trans. Amer. Math. Soc. {\bf 74}, (1953). 110–134.  
\bibitem{Ksy} H. Kasuya, Cohomologically symplectic solvmanifolds are symplectic. J. Symplectic Geom.  {\bf 9} (2011), no. 4, 429--434. 
\bibitem{Kfo} H. Kasuya, Formality and hard Lefschetz property of aspherical manifolds. Osaka J. Math. {\bf 50} (2013), no. 2, 439--455.
\bibitem{Kas}
H. Kasuya, Minimal models, formality and Hard Lefschetz properties  of solvmanifolds with local systems, J. Differential Geom., {\bf 93}, (2013), 269--298.

\bibitem{KCT} H. Kasuya,
Central theorems for cohomologies of certain solvable groups.
 arXiv:1311.1310 to appear in Transactions of the AMS.
\bibitem{KSS} H. Kasuya, Cohomologies of Sasakian groups and Sasakian solvmanifolds, arXiv:1412.5825 To appear in Ann. Mat. Pura Appl
\bibitem{Jan}
J. C. Jantzen,  Representations of algebraic groups. Pure and Applied Mathematics, {\bf 131}. Academic Press, Inc., Boston, MA, 1987.
\bibitem{LP} L. Lambe, S. Priddy, Cohomology of nilmanifolds and torsion-free, nilpotent groups, Trans. Amer. Math. Soc {\bf 273} (1982) 39–-55. 
\bibitem{LS} P. Lambrechts, D.  Stanley
The rational homotopy type of a blow-up in the stable case. Geom. Topol. {\bf 12} (2008), no. 4, 1921–1993.
%\bibitem{LM} A. Lubotzky, A. R Magid, Varieties of representations of finitely generated groups. Mem. Amer. Math. Soc. 58 (1985), no. 336

\bibitem{Li}
Y. Lin, Lefschetz contact manifolds and odd dimensional symplectic geometry, arXiv:1311.1431v2 (2013).
%\bibitem{Mc}J. McCleary,  A user's guide to spectral sequences, Second edition, Cambridge Studies in Advanced Mathematics, \textbf{58}, Cambridge University Press, Cambridge, 2001.
\bibitem{Mcd} D. McDuff
Examples of simply-connected symplectic non-K\"ahlerian manifolds. J. Differential Geom. {\bf 20} (1984), no. 1, 267--277. 
%\bibitem{Mor} J. W. Morgan, The algebraic topology of smooth algebraic varieties. Inst. Hautes Études Sci. Publ. Math. No. {\bf 48} (1978), 137–204.
\bibitem{Mos1} G. D. Mostow, Fully reducible subgroups of algebraic groups. Amer. J. Math. {\bf 78} (1956), 200--221.

%\bibitem{Mosc} G. D. Mostow,  Cohomology of topological groups and solvmanifolds. Ann. of Math. (2) {\bf 73} 1961 20--48.

%\bibitem{Na} I. Nakamura, Complex parallelisable manifolds and their small deformations. J. Differential Geometry {\bf 10} (1975), 85--112.
%\bibitem{Nar}A. Narkawicz, Cohomology Jumping Loci and Relative Malcev Completion.
%http://arxiv.org/abs/0804.4164
\bibitem{Nom}
K. Nomizu, On the cohomology of compact homogeneous spaces of nilpotent Lie groups. 
Ann. of Math. (2) {\bf 59}, (1954). 531--538.
%\bibitem{OV2}A. L. Onishchik, and E. B. Vinberg, Structure of Lie groups and Lie algebras, Springer (1994).
%\bibitem{OV}A. L. Onishchik, E. B. Vinberg, Discrete subgroups of Lie groups and cohomologies of Lie groups and Lie algebras, Springer(2000).
%\bibitem{OT}K. Oeljeklaus, M. Toma,  Non-K\"ahler compact complex manifolds associated to number fields. Ann. Inst. Fourier (Grenoble) {\bf 55} (2005), no. 1, 161--171.
 \bibitem{R}
 M. S. Raghunathan, Discrete subgroups of Lie Groups, Springer-Verlag, New York, 1972.
%   \bibitem{Re} B. E. Reed,  Representations of solvable Lie algebras, Michigan Math. J.  {\bf 16} 1969 227--233. 
%\bibitem{Saw} H. Sawai, A construction of lattices on certain solvable Lie groups. Topology Appl. {\bf 154} (2007), no. 18, 3125--3134.
%\bibitem{Sim} C. T. Simpson, Higgs bundles and local systems. Inst. Hautes $ \rm \acute E$tudes Sci. Publ. Math. No. {\bf 75} (1992), 5--95. 
\bibitem{Shul}H. Shulman, D. Tischler,
Leaf invariants for foliations and the Van Est isomorphism. 
J. Differential Geometry {\bf 11} (1976), no. 4, 535--546.

\bibitem{Sul} D. Sullivan, 
Infinitesimal computations in topology. Inst. Hautes $ \rm \acute E$tudes Sci. Publ. Math. No. {\bf 47} (1977), 269--331 (1978).
\bibitem{Ts}
D. Tischler, Closed $2$–forms and an embedding theorem for manifolds, J. Differential Geometry {\bf 12} (1977), 229--235.
\end{thebibliography}

\documentclass[11pt]{amsart}
\title[Extended  simplicial rational Nomizu's Theorem]
{Extended  simplicial rational Nomizu's Theorem}

\author{Hisashi Kasuya}

\usepackage{amssymb}
\usepackage{amsmath}
\usepackage{amscd}
\usepackage{amstext}
\usepackage{amsfonts}
\usepackage[all]{xy}

\theoremstyle{plain}

\theoremstyle{plain}

\theoremstyle{plain}

\theoremstyle{plain}
\newtheorem{theorem}{Theorem}[section] 
\theoremstyle{remark}
\newtheorem{remark}{Remark}[section] 
\theoremstyle{remark}
\newtheorem{Important note}[theorem]{Important note}
\theoremstyle{Main result}
\newtheorem{main result}{Main result}
\theoremstyle{lemma}

\theoremstyle{definition}

\theoremstyle{proposition}
\newtheorem{proposition}[theorem]{Proposition}
\theoremstyle{corollary}
\newtheorem{corollary}[theorem]{Corollary}
\theoremstyle{remark}
\newtheorem{example}{Example}
\theoremstyle{plain}
\newtheorem{newexample}{New Examples}

\address[Hisashi Kasuya]{Department of Mathematics, Tokyo Institute of Technology, 1-12-1, O-okayama, Meguro, Tokyo 152-8551, Japan}
\email{kasuya@math.titech.ac.jp}

\keywords{Sullivan's minimal model, polycyclic group,
rational cohomology of algebraic group, simplicial classifying space, simplicial de Rham theory, Borel construction}
\subjclass[2010]{Primary:20F16, 20G15, 55R35, 55P62  Secondary:20G10, 17B45, 55U10}

\newcommand{\C}{\mathbb{C}}
\newcommand{\R}{\mathbb{R}}

\newcommand{\Q}{\mathbb{Q}}
\newcommand{\Z}{\mathbb{Z}}

\newcommand{\n}{\frak{n}}

\begin{document} 

\maketitle
\begin{abstract}

For a torsion-free virtually polycyclic group $\Gamma$,
we give a canonical homomorphism form  certain finite-dimensional cochain complex to the $\Q$-polynomial de Rham complex of the simplicial classifying space $B\Gamma$ which induces a cohomology isomorphism.
By this result, we obtain the Sullivan's  minimal model of certain differential graded algebra  defined on $B\Gamma$ and 
we obtain new examples of hard Lefschetz  symplectic manifolds and hard Lefschetz  contact manifolds.
\end{abstract}

\section{Introduction}

A group $\Gamma$ is polycyclic if it admits a sequence 
\[\Gamma=\Gamma_{0}\supset \Gamma_{1}\supset \cdot \cdot \cdot \supset \Gamma_{k}=\{ e \}\]
of subgroups such that each $\Gamma_{i}$ is normal in $\Gamma_{i-1}$ and $\Gamma_{i-1}/\Gamma_{i}$ is cyclic.
For a polycyclic group $\Gamma$, we denote ${\rm rank}\,\Gamma=\sum_{i=1}^{i=k} {\rm rank}\,\Gamma_{i-1}/\Gamma_{i}$.
Let $\Gamma$ be a torsion-free virtually polycyclic group.
For a representation of $\Gamma$ into a $\Q$-algebraic group $G$,
 if the image $\rho(\Gamma)$ is Zariski-dense in $G$, then  we have $\dim U\le {\rm rank}\,\Gamma$ where $U$ is the  unipotent radical of $G$  (see \cite[Lemma 4.36.]{R}).
We say that  $\rho:\Gamma\to  G$ is a full representation if $\dim U= {\rm rank}\,\Gamma$.

For a  a simplicial complex $K$ with $\pi_{1}K=\Gamma$ and a $\Gamma$-module $V$ with a finite dimensional $\Q$-vector space, considering $V$ as a local system on $K$,
we can define  the $\Q$-polynomial de Rham complex  $A^{\ast}_{p}(K, V)$ of $K$ with values in local system $V$.
We also consider $V=\varinjlim V_{i}$ for an inductive system of finite dimensional $\pi_{1}(K)$-modules $V_{i}$.
We define
\[A^{\ast}_{p}(K, V)=\varinjlim A^{\ast}_{p}(K, V_{i}).
\]

Let $\Gamma$ be a torsion-free virtually polycyclic group.
We consider the following situation:
\begin{itemize}
\item We have a $\Q$-algebraic group $G$ and an injective representation $\rho :\Gamma\to G$ such that the image $\rho(\Gamma)$ is Zariski-dense in $G$.
\item $\rho:\Gamma\to  G$ is a full representation.
\end{itemize}
We consider the simplicial classifying space $B\Gamma$ of a torsion-free virtually polycyclic group $\Gamma$.
For a rational  $G$-module $V$,
we consider the $\Q$-polynomial de Rham complex  $A^{\ast}_{p}(B\Gamma, V)$.
We also consider the complex of "$G$-invariant differential forms" on $U$.
We can take
a splitting $G=T\ltimes U$ such that $U$ is the unipotent radical of $G$ and $T$ is a maximal reductive subgroup of $G$ (see \cite{Mos1}).
For the Lie algebra $\frak u$ of $U$, we consider the cochain complex
$\left(\bigwedge {\frak u}^{\ast}\otimes V\right)^{T}$  of the $T$-invariant elements of the cochain complex of the Lie algebra ${\frak u}$ with values in $V$.
Then, in this paper we show the following result.
\begin{theorem}\label{MINt}
We have an explicit  map $\left(\bigwedge {\frak u}^{\ast}\otimes V\right)^{T}\to  A^{\ast}_{p}(B\Gamma,V)$ which induces a cohomology isomorphism.
\end{theorem}

\begin{remark}
It is known that a $\Q$-algebraic group $G$ and an injective representation $\rho :\Gamma\to G$  so that
\begin{itemize}
\item We have a $\Q$-algebraic group $G$ and an injective representation $\rho :\Gamma\to G$ such that the image $\rho(\Gamma)$ is Zariski-dense in $G$.
\item $\rho:\Gamma\to  G$ is a full representation.
\item  The centralizer $Z_{G}(U)$ of $U$ is contained in $U$.
\end{itemize}
 exist and such $G$ is unique up to isomorphism of $\Q$-algebraic groups (\cite[Appendix A.]{B}).
Such $G$ is called the algebraic hull of $\Gamma$.

For a representation $\phi:\Gamma\to GL(V)$ with a finite dimensional $\Q$-vector space $V$,
consider the representation $\rho\times \phi:\Gamma\to G\times GL(V)$ and the take the Zariski-closure $G^{\phi}$ of the its image.
Then the representation $\Gamma\to G^{\phi}$ is also a full representation and $\phi:\Gamma\to GL(V)$ is extended to  a rational representation $G^{\phi}\to GL(V)$ (see \cite[Section 3]{KCT}).
Thus, for any representation $\phi:\Gamma\to GL(V)$ with a finite dimensional $\Q$-vector space $V$, applying Theorem \ref{MINt}, the cohomology of the $\Q$-polynomial de Rham complex  $A^{\ast}_{p}(B\Gamma, V)$ is computed by the finite-dimensional cochain complex $\left(\bigwedge {\frak u}^{\ast}\otimes V\right)^{T}$.
\end{remark}

Theorem \ref{MINt} can be regarded as a generalization of  simplicial rational version of Nomizu's Theorem (\cite{Nom}, \cite{LP}).
Let $N$ be a simply connected nilpotent Lie group and $\frak n$ be the Lie algebra of $N$.
We suppose that  $N$ has a lattice (i.e. cocompact discrete subgroup) $\Gamma$.
We consider the nilmanifold $\Gamma\backslash N$.
Then, considering the cochian complex $\bigwedge \frak n^{\ast}$ of the Lie algebra as the space of the invariant differential forms, in \cite{Nom}, Nomizu proves that
the canonical inclusion $\bigwedge \frak n^{\ast}\subset A^{\ast}(\Gamma\backslash N)$ induces a cohomology isomorphism
\[H^{\ast}(\n,\R)\cong H^{\ast}(\Gamma\backslash N,\R)
\]
where $A^{\ast}(\Gamma\backslash N)$ is the de Rham complex of $\Gamma\backslash N$.

In \cite{LP}, Lambe and Priddy give a  simplicial rational version of  Nomizu's theorem.
For simply connected nilpotent Lie group $N$ with a lattice $\Gamma$, $\Gamma$ is a torsion-free finitely generated nilpotent group and a nilmanifold $\Gamma\backslash N$ is an aspherical manifold with the fundamental group $\Gamma$.
Conversely any torsion-free finitely generated nilpotent group $\Gamma$ can be embedded in a simply connected nilpotent Lie group $N$ whose Lie algebra $\frak n$ admits a $\Q$-structure $\frak n_{\Q}$ (see \cite{R}).
In \cite{LP}, considering the simplicial classifying space $B\Gamma$ of a torsion-free finitely generated nilpotent group $\Gamma$ and the $\Q$-polynomial de Rham complex  $A^{\ast}_{p}(B\Gamma,\Q)$,  Lambe and Priddy construct an explicit map $\bigwedge \frak n_{\Q}\to A^{\ast}_{p}(B\Gamma,\Q)$ which induces a cohomology isomorphism.
A simply connected nilpotent Lie group $N$ can be considered as a real unipotent algebraic group with a $\Q$-structure  $N(\Q)$ so that $\Gamma\subset N(\Q)$.
Regarding a torsion-free finitely generated nilpotent group $\Gamma$ as a polycyclic group, we can say that the $\Q$-algebraic group $N(\Q)$ is the algebraic hull of $\Gamma$ (see \cite{R}).

Nomizu's theorem gives an important fact on the theory of Sullivan's minimal model.
We can say that the Differential graded algebra (shortly DGA) $\bigwedge \frak n^{\ast}$ is the  minimal model of  $A^{\ast}(\Gamma\backslash N)$ (see \cite{H}).

We can generalize this fact.
For a  torsion-free virtually polycyclic group $\Gamma$,
take $G$ the algebraic hull of $\Gamma$ and a splitting $G=T\ltimes U$ for a maximal reductive subgroup $T$.
Denote by $\frak u$  the Lie algebra of  the unipotent hull $U$ of $\Gamma$.
Consider the the coordinate ring $\Q[T]$ of the algebraic group $T$.
By the composition $\Gamma\to G\to T$ where $G\to T$ is the projection,
since $\Q[T]$ is a rational $T$-module, we regard $\Q[T]$ as a $\Gamma$-module.
Consider the $\Q$-polynomial de Rham complex  $A^{\ast}_{p}(B\Gamma,\Q[T])$.
Then $A^{\ast}_{p}(B\Gamma,\Q[T])$ is a differential graded algebra (DGA).
By Theorem \ref{MINt}, we have the following result.

\begin{theorem}\label{polmi}
We have an explicit DGA map $\bigwedge {\frak u}^{\ast}\to  A^{\ast}_{p}(B\Gamma,\Q[T])$ which induces a cohomology isomorphism.
Hence $\bigwedge{\frak u}^{\ast}$ is the minimal model of $A^{\ast}_{p}(B\Gamma,\Q[T])$.
\end{theorem}

See \cite{Kas} for the similar result on solvmanifolds where a solvmanifold is the compact  quotient $\Gamma\backslash S$
of a simply connected solvable Lie group $S$ by a lattice $\Gamma$.

An infra-solvmanifold is a manifold of the form $G/\Delta$, where $G$ is a simply connected solvable Lie group, and $\Delta$ is a torsion-free subgroup of ${\rm Aut}(G)\ltimes G$ such that the closure of $h(\Delta)$ in ${\rm Aut}(G)$ is compact where $h:{\rm Aut}(G)\ltimes G\to {\rm Aut}(G)$ is the projection.
An infra-solvmanifold is a  generalization of a solvmanifold.
An infra-solvmanifold is a aspherical manifold with a virtually polycyclic group.
Theorem \ref{MINt} is useful for finding symplectic structures on infra-solvmanifolds $M$ whose cohomology classes in $H^{2}(M,\Z)$.
By finding such symplectic infra-solvmanifolds,
we obtain the following results.
\begin{theorem}
We obtain the following new examples.
\begin{itemize}
\item Symplectic blow-ups of   complex projective spaces along 
certain embedded infra-solvmanifolds which satisfy the hard Lefschetz properties.
(But, it is not clear whether these manifolds admit K\"ahler metrics.)

\item Non Sasakian contact manifolds which satisfy the hard Lefschetz properties in the sense of \cite{CNY}, \cite{Li}.

\end{itemize}

\end{theorem}

\section{Cohomology of algebraic groups}\label{SAL}
The purpose of this section is to construct an explicit cochain complex homomorphism which induces the  Hochschild isomorphism on the rational cohomology of an algebraic group.
For this construction, we are inspired by the simplicial construction of Van Est isomorphism on the continuous cohomology of a Lie group (see \cite{Shul}). 
Let $G$ be a $\Q$-algebraic group.
A  $G$-module is called rational if it is  the sum of finite-dimensional $G$-stable subspaces $\{V_{i}\}$ so that each $V_{i}$ comes from a rational representation. 
For a rational  $ G$-module $V$, we define the rational cohomology $H^{\ast}( G,V)={\rm Ext}^{\ast}_{ G}(\Q,V)$ as \cite{Hoc} and \cite{Jan}.
We have the standard resolution. 
We denote by $C^{p}( G,V)$ the set of the $V$-valued rational functions on $\underbrace{G\times \dots\times G }_{p+1}$ with the left-$ G$-action.
Consider the sequence
\[\xymatrix{
V\ar[r]& C^{0}(G,V)\ar^{d}[r]&C^{1}( G,V)\ar^{d}[r]&\dots
 }
\] 
such that the first map $V\to C^{0}( G,V)$ is the embedding as  constant functions and $d:C^{p}( G,V)\to C^{p+1}( G,V)$ is given by
\[d\phi (g_{0},\dots,g_{p+1})=\sum (-1)^{i}\phi(g_{0},\dots, {\hat g}_{i},\dots, g_{p+1})
\]
for $\phi\in C^{p}( G,V)$, $g_{0},\dots, g_{p+1}\in G$.
Then the rational cohomology $H^{\ast}(G,V)$ is the cohomology of the cochain complex $C^{\ast}( G,V)^{ G}$.
For the unipotent radical $U$ of $G$ and a maximal reductive subgroup $T$, we have a splitting $G=T\ltimes U$ (\cite{Mos1}).
Let $\frak u$ be the Lie algebra of $U$.
Hochschild showed that we have an isomorphism
\[H^{\ast}({\frak u}, V)^{T}\cong H^{\ast}(G,V).
\]
The purpose of this section is to represent this isomorphism 
 as a cochain complex homomorphism.

Let $A^{\ast}_{a}(U)$ be the algebraic de Rham complex of the algebraic variety $U$.
Consider the coordinate ring $\Q[U]$ of $U$.
By the spectral sequence as \cite[Proposition 3.4]{Haia}, we have $H^{1}(G,\Q[U] \otimes W)=H^{1}(U,\Q[U] \otimes W)^T=0$ for any finite dimensional rational $G$-module $W$.
This implies that $\Q[U]$   is an injective $G$-module (see \cite{Hoi})
We have $A^{\ast}_{a}(U)\cong {\rm Hom}_{\Q}(\bigwedge {\frak u}, \Q[U])$  and hence $A^{\ast}_{p}(U)\otimes V$ is an injective $G$-module (see \cite{Jan}).
Since the exponential map $\exp: {\frak u}\to U$ is an isomorphism of $\Q$-algebraic variety,  we have $H^{0}(A^{\ast}_{a}(U)\otimes V)=V$ and $H^{\ast}(A^{\ast}_{a}(U)\otimes V)=0$ for $\ast>0$.
Hence the sequence 
 \[\xymatrix{
V\ar[r]& A_{a}^{0}(U)\otimes V\ar^{d}[r]&A_{a}^{1}(U)\otimes V\ar^{d}[r]&\dots
 }
\] 
is an injective resolution.

By a splitting $G=T\ltimes U$, we have the homomorphism $\alpha : G\to {\rm Aut}(U)\ltimes U$.
We define the map 
\[\sigma^{p}(\cdot)(\cdot):\underbrace{G\times \dots\times G }_{p+1}\times \Q^{p}\to U\]
 such that
$\sigma ^{0}(g_{0})(0)=\alpha(g_{0})e$, $\sigma^{1}(g_{0},g_{1})(t_{1})=\alpha(g_{0}){\rm exp}((1-t_{1}){\rm log} \alpha(g_{0}^{-1}g_{1})e$ and inductively 
\[\sigma^{p}(g_{0},\dots, g_{p})(t_{1},\dots, t_{p})=\alpha(g_{0}){\rm exp}((1-t_{1}){\rm log} \sigma^{p-1}(g_{0}^{-1}g_{1},\dots, g_{0}^{-1}g_{p})(t_{2},\dots, t_{p})).
\]
It is known that the exponential map ${\rm exp}:\frak u\to U$ is an isomorphism of $\Q$-algebraic variety and ${\rm log} :U\to \frak u$ is the inverse.
Hence  the map 
\[\sigma^{p}:\underbrace{G\times \dots\times G }_{p+1}\times \Q^{p}\to U\]
 is a  homomorphism of $\Q$-algebraic variety such that 
for any $(t_{1},\dots,t_{p})$, the map 
\[\sigma^{p}(\cdot)(t_{1},\dots,t_{p}):\underbrace{G\times \dots\times G }_{p+1} \to U\]
 is a $G$-equivariant map.
We note 
\begin{multline*}
\sigma^{p}(g_{0},\dots,g_{p})(t_{1},\dots t_{i-1},0,t_{i},\dots t_{p-1})\\
=\sigma^{p-1}(g_{0},\dots ,g_{i-1},g_{i+1},\dots ,g_{p})(t_{1},\dots, t_{p-1}).
\end{multline*}

 For  $(g_{0},\dots,g_{p})$  and $\omega\in A^{\ast}_{a}(U)\otimes V$, considering the map $\sigma^{p}(g_{0},\dots,g_{p})(\cdot):\Q^{p} \to U\cong {\frak u}$ which is a homomorphism of  $\Q$-algebraic variety, we have the $\Q$-algebraic differential form $\sigma^{p}(g_{0},\dots,g_{p})^{\ast}\omega$ for parameters $(t_{1},\dots,t_{p})\in \Q^{p}$.
We regard $\sigma^{p}(g_{0},\dots,g_{p})^{\ast}\omega$ as a $\Q$-polynomial differential form on $\R^{p}$.
We define the map $\theta : A^{p}_{a}(U)\otimes V\to C^{p}(G,V)$ such that 
\[\theta (\omega)(g_{0},\dots,g_{p})=\int_{\Delta}\otimes {\rm id}_{V}\sigma^{p}(g_{0},\dots,g_{p})^{\ast}\omega\]
 where 
\[\Delta=\left\{(1-t_{1}-\dots-t_{p},t_{1},\dots,t_{p})\vert 0\le t_{i}\le 1  \right\}
\] is the standard $p$-simplex for the parameters $(t_{1},\dots,t_{p})$.
Then we can easily show that 
the map $\theta :   A^{\ast}_{a}(U)\otimes V\to C^{\ast}(G, V)$ is $G$-equivariant cochain complex homomorphism
by  Stokes' theorem. (cf.  \cite[Section 3]{Shul})

Now we have $(A^{\ast}_{a}(U)\otimes V)^{G}=\left(\bigwedge {\frak u}^{\ast}\otimes V\right)^{T}$
 where $\bigwedge {\frak u}^{\ast}\otimes V$ is the cochain complex of the Lie algebra ${\frak u}$ with values in the $\frak u$-module $V$.
Consider the restriction $\theta: \left(\bigwedge {\frak u}^{\ast}\otimes V\right)^{T}\to C^{\ast}(G, V)^{G}$.
Then the induced map $\theta:H^{\ast}({\frak u}, V)^{T}\to H^{\ast}(G,V)$ is identified with the map  ${\rm Ext}^{\ast}_{\mathcal G}(\Q,V)\to {\rm Ext}^{\ast}_{\mathcal G}(\Q,V)$ induced by the identity map $V\to V$.
Hence we have the following result.
\begin{theorem}\label{ALGCO}
The map $\theta: \left(\bigwedge {\frak u}^{\ast}\otimes V\right)^{T}\to C^{\ast}(G, V)^{G}$ induces a cohomology isomorphism
\[H^{\ast}({\frak u}, V)^{T}\cong H^{\ast}(G,V).
\]
\end{theorem}

\section{Simplicial  de Rham theory}\label{des}
We denote by $A^{\ast}_{p}(n)$ the $\Q$-DGA which is generated by $t_{0},\dots, t_{n}$ of degree $0$ and $dt_{0},\dots, dt_{n}$ of degree $1$ with the relations $t_{0}+\dots+t_{n}=1$ and $dt_{0}+\dots+dt_{n}=0$.
We can regard  $A^{\ast}_{p}(n)$ as the  $\Q$-polynomial  de Rham complex on the standard $n$-simplex $\Delta^{n}$.
We define the map $\int_{\Delta^{n}} : A^{n}_{p}(n)\to \Q$ by the ordinary Riemannian integral.
Let $K$ be a simplicial complex with a universal covering complex $\tilde K$.
We denote by $n(\sigma)$ the dimension of a simplex $\sigma\in K$.
Let $V$ be a finite-dimensional $\Q$-vector space which is a $\pi_{1}(K)$-module.
We denote by  $A^{\ast}_{p}( K, V)$ the space of collections
$\{\omega_{\sigma}\in A^{\ast}_{p}(n(\sigma))\otimes V \}_{\sigma\in \tilde K}$ such that:
\begin{itemize}
\item $\{\omega_{\sigma}\}_{\sigma\in \tilde K}$ are  compatible under restrictions  to faces i.e. $i^{\ast}\omega_{\sigma}=\omega_{\tau}$ for the inclusion $i:\tau\to \sigma$ of a face.
\item $\{\omega_{\sigma}\}_{\sigma\in \tilde K}$ are invariant under the $\pi_{1}(K)$-action i.e.
$\gamma\cdot \omega_{\gamma\sigma}=\omega_{\sigma}$ for any $\gamma\in \pi_{1}(K)$.
\end{itemize}
We call  $A^{\ast}_{p}(K, V)$ the $\Q$-polynomial de Rham complex of $K$ with values in local system $V$.
The space $A^{\ast}_{p}(K, V)$ with the exterior derivation is a cochain complex.
Let $C^{\ast}(K,V)=(C^{\ast}(\tilde K)\otimes V)^{\Gamma}$ be the cochain complex of simplicial cochains with values in the local system $V$.
Define the map $\iota: A^{n}_{p}( K, V)\to C^{n}(K,V)$ such that for $\sigma\in K$ with $n(\sigma)=n$
\[\iota(\{\omega\})(\sigma)=\int_{\Delta^{n}}\otimes {\rm id}_{V} (\omega_{\sigma}).
\]
Then, 
this map is a cochain complex homomorphism and this map induces a cohomology isomorphism
(see \cite[Chapter 9]{GMo},  \cite[Chapter 12-14]{Halp}).

Let $V=\varinjlim V_{i}$ for an inductive system of finite dimensional $\pi_{1}(K)$-modules.
We define
\[A^{\ast}_{p}(K, V)=\varinjlim A^{\ast}_{p}(K, V_{i})
\]

\begin{example}\label{HainDG}{\rm(cf.\cite{Hai}) }
Let $T$ be a reductive $\Q$-algebraic group and $\rho:\pi_{1}(K)\to T$ a representation.
Consider the coordinate ring $\Q[T]$   of $T$.
Then as a $(T,T)$-bimodule,
we have 
\[\Q[T]=\bigoplus V_{\alpha}^{\ast}\otimes V_{\alpha}\]
such that $\{V_{\alpha}\}$ is a set of isomorphism classes of irreducible right $T$-module (\cite[Proposition 3.1]{Hai}).
We regard $\Q[T]$ as a $\pi_{1}(K)$-module by $\rho$.
Then we have 
\[A^{\ast}_{p}(K, \Q[T])=\bigoplus A^{\ast}_{p}(K, V_{\alpha}^{\ast})\otimes V_{\alpha}
\] 
and it is a DGA.

\end{example}


Let $\Gamma$ be a discrete group.
For a $\Gamma$-module $V$, we define the group cohomology $H^{\ast}(\Gamma,V)={\rm Ext}^{\ast}_{\Gamma}(\Q,V)$.
We have the standard resolution.
We denote by $C^{p}(\Gamma,V)$ the set of the $V$-valued functions on $\underbrace{\Gamma\times \dots\times \Gamma }_{p+1}$ with the left-$\Gamma$-action.
Consider the sequence
\[\xymatrix{
V\ar[r]& C^{0}(\Gamma,V)\ar^{d}[r]&C^{1}(\Gamma,V)\ar^{d}[r]&\dots
 }
\] 
such that the first map $V\to C^{0}(\Gamma,V)$ is the embedding as  constant functions and $d:C^{p}(\Gamma,V)\to C^{p+1}(\Gamma,V)$ is given by
\[d\phi (\gamma_{0},\dots,\gamma_{p+1})=\sum (-1)^{i}\phi(\gamma_{0},\dots, {\hat \gamma}_{i},\dots, \gamma_{p+1})
\]
for $\phi\in C^{p}(\Gamma,V)$, $\gamma_{0},\dots, \gamma_{p+1}\in\Gamma$.
Then the group cohomology $H^{\ast}(\Gamma,V)$ is the cohomology of the cochain complex $C^{\ast}(\Gamma,V)^{\Gamma}$.
A representation $\rho:\Gamma\to G$ induces a homomorphism $\rho^{\ast}:H^{\ast}(G,V)\to H^{\ast}(\Gamma,V)$.
By the map $\theta: \left(\bigwedge {\frak u}^{\ast}\otimes V\right)^{T}\to C^{\ast}(G, V)^{G}$ as Section \ref{SAL},
we give a geometric representation of $\rho^{\ast}:H^{\ast}(G,V)\to H^{\ast}(\Gamma,V)$.

We define the acyclic simplicial complex $E\Gamma$ with the free discontinuous $\Gamma$-action so that:
\begin{itemize}
\item For integers $n\ge 0$, simplices of $E\Gamma$ are standard $n$-simplices $\Delta_{(\gamma_{0},\dots,\gamma_{n})}$ indexed by $\Gamma^{n+1}$.
\item \[(t_{1},\dots,t_{i-1},t_{i+1},\dots,t_{n})_{(\gamma_{0},\dots,\gamma_{i-1},\gamma_{i+1},\dots,\gamma_{n})}\in \Delta_{(\gamma_{0},\dots,\gamma_{i-1},\gamma_{i+1},\dots,\gamma_{n})}\] is identified with
 \[(t_{1},\dots,t_{i-1},0,t_{i+1},\dots,t_{n})_{(\gamma_{0},\dots,\gamma_{n})}\in \Delta_{(\gamma_{0},\dots,\gamma_{n})}.\]
\item For $\gamma\in\Gamma$, the action is given by
\[\gamma\cdot (t_{1},\dots,t_{n})_{(\gamma_{0},\dots,\gamma_{n})}=(t_{1},\dots,t_{n})_{(\gamma\gamma_{0},\dots,\gamma\gamma_{n})}.
\]

\end{itemize}
We define  $B\Gamma$ the quotient of $E\Gamma$ by the $\Gamma$-action.
Then the simplicial complex $B\Gamma$ is an Eilenberg-Maclane space $K(\Gamma,1)$.
For a finite dimensional $\Gamma$-module $V$, we consider  the $\Q$-polynomial de Rham complex $A^{\ast}_{p}(B\Gamma,V)$ of $B\Gamma$ with values in local system $V$.
Define the map $\iota:  A^{n}_{p}(B\Gamma,V)\to C^{n}(\Gamma,V)^{\Gamma}$ as 
\[\iota(\{\omega_{\sigma}\}_{\sigma\in B\Gamma} )(\gamma_{0},\dots,\gamma_{n})=\int_{\Delta_{(\gamma_{0},\dots,\gamma_{n})}}\omega_{\Delta_{(\gamma_{0},\dots,\gamma_{n})}}.
\]
Since we can identify $C^{\ast}(\Gamma,V)^{\Gamma}$ with the cochain complex  $C^{\ast}(B\Gamma,V)$, the map $\iota:  A^{p}_{p}(B\Gamma,V)\to C^{p}(\Gamma,V)^{\Gamma}$ induces a cohomology isomorphism
\[H^{\ast}(A^{\ast}_{p}(B\Gamma,V))\cong H^{\ast}(\Gamma,V).
\]
We define the map $\psi:\left(\bigwedge {\frak u}^{\ast}\otimes V\right)^{T}\to  A^{\ast}_{p}(B\Gamma,V)$ such that 
\[\psi(\omega)=\{\sigma^{p}(\rho(\gamma_{0}),\dots,\rho(\gamma_{p}))^{\ast}\omega\}_{\Delta_{(\gamma_{0},\dots,\gamma_{n})}}\]
where $\sigma^{p}(\rho(\gamma_{0}),\dots,\rho(\gamma_{p}))^{\ast}\omega$ is defined in Section \ref{SAL}.
By the $G$-invariance of $\omega\in \left(\bigwedge {\frak u}^{\ast}\otimes V\right)^{T}$   and the relation
\begin{multline*}\sigma^{p}(g_{0},\dots,g_{p})(t_{1},\dots t_{i-1},0,t_{i},\dots t_{p-1})
\\
=\sigma^{p-1}(g_{0},\dots ,g_{i-1},g_{i+1},\dots ,g_{p})(t_{1},\dots, t_{p-1}),
\end{multline*}
we actually have $\psi(\omega)\in  A^{\ast}_{p}(B\Gamma,V)$.
Then 
the map $\psi:\left(\bigwedge {\frak u}^{\ast}\otimes V\right)^{T}\to  A^{\ast}_{p}(B\Gamma,V)$ is a cochain complex homomorphism.

We have the commutative diagram
\[\xymatrix{
\left(\bigwedge {\frak u}^{\ast}\otimes V\right)^{T}\ar[d]^{\psi}\ar[r]^{\theta}&C^{\ast}( G,V)^{G}\ar[d]^{\rho^{\ast}}\\
A^{\ast}_{p}(B\Gamma,V)\ar[r]^{\iota}& C^{p}(\Gamma,V)^{\Gamma}.
}
\]
Hence we have:
\begin{corollary}\label{cocBB}
The induced map $\psi^{\ast}:H^{\ast}({\frak u}, V)^{T}\to  H^{\ast}(A^{\ast}_{p}(B\Gamma,V))$ is identified with the map $\rho^{\ast}:H^{\ast}(G,V)\to H^{\ast}(\Gamma,V)$.
\end{corollary}
Suppose $V=\Q[T]$.
Then we have $\left(\bigwedge {\frak u}^{\ast}\otimes \Q[T]\right)^{T}=\bigwedge {\frak u}^{\ast}$ and we have the DGA map $\psi:\bigwedge {\frak u}^{\ast}\to A^{\ast}_{p}(B\Gamma,\Q[T])$.
As we notice in Example \ref{HainDG},
in this case Corollary \ref{cocBB} is written as the following statement.
\begin{corollary}\label{Hni}
The  induced map $\psi^{\ast}:H^{\ast}({\frak u}, \Q)\to  H^{\ast}(A^{\ast}_{p}(B\Gamma,\Q[T]))$ is identified with the map \[\rho^{\ast}:\bigoplus H^{\ast}(G,V_{\alpha}^{\ast})\otimes V_{\alpha}\to \bigoplus H^{\ast}(\Gamma,V_{\alpha}^{\ast})\otimes V_{\alpha}.\]

\end{corollary}

\section{Classifying spaces of torsion-free virtually polycyclic groups}\label{poll}
Let $\Gamma$ be a torsion-free virtually polycyclic group,
  $G$ a $\Q$-algebraic group and  $\rho:\Gamma\to  G$ a representation with the Zariski-dense image.
It is known that we have $\dim U\le {\rm rank}\,\Gamma$ where $U$ is the  unipotent radical of $G$.
We say that  $\rho:\Gamma\to  G$ is a full representation if $\dim U= {\rm rank}\,\Gamma$.
\begin{theorem}[\cite{KCT}]\label{kasct}
If $\rho:\Gamma\to  G$ is an injective full representation, then for any rational $G$-module $V$ the induced map 
 $\rho^{\ast}:H^{\ast}(G,V)\to H^{\ast}(\Gamma,V)$
is an isomorphism.
\end{theorem}
We suppose that $\rho:\Gamma\to  G$ is an injective full representation.
Denote by $\frak u$  the Lie algebra of  the unipotent radical  $U$ of $G$.
Take a splitting $G=T\ltimes U$ for a maximal reductive subgroup $T$.
Then by Corollary \ref{cocBB} and Theorem \ref{kasct}, we have the following fact.

\begin{theorem}\label{indisoinn}
 The map $\psi:\left(\bigwedge {\frak u}^{\ast}\otimes V\right)^{T}\to  A^{\ast}_{p}(B\Gamma,V)$ induces a cohomology isomorphism.
\end{theorem}
Suppose  $V=\Q[T]$ as a $\Gamma$-module.
By Corollary \ref{Hni}, we obtain the following result.

\begin{theorem}\label{polmi}
We have an explicit DGA map $\bigwedge {\frak u}^{\ast}\to  A^{\ast}_{p}(B\Gamma,\Q[T])$ which induces a cohomology isomorphism.
Hence $\bigwedge{\frak u}^{\ast}$ is the minimal model of $A^{\ast}_{p}(B\Gamma,\Q[T])$.
\end{theorem}

\section{Examples and applications to symplectic and contact geometry}\label{Exss}

 An infra-solvmanifold is a manifold of the form $G/\Delta$, where $G$ is a simply connected solvable Lie group, and $\Delta$ is a torsion-free subgroup of ${\rm Aut}(G)\ltimes G$ such that the closure of $h(\Delta)$ in ${\rm Aut}(G)$ is compact where $h:{\rm Aut}(G)\ltimes G\to {\rm Aut}(G)$ is the projection.
An infra-solvmanifold is a  generalization of a solvmanifold.
An infra-solvmanifold is a aspherical manifold with a virtually polycyclic group.

Let $\Gamma$ be a torsion-free virtually polycyclic group.
Then there exists a compact infra-solvmanifold $M$ with the fundamental group $\Gamma$ (see \cite{B}).
It is known that every compact infra-solvmanifold is smoothly rigid   (\cite[Corollary 1.5]{B}).
Hence we have the canonical correspondence between   torsion-free virtually polycyclic groups $\Gamma$ and  infra-solvmanifolds $M_{\Gamma}$ with $\pi_{1}(M_{\Gamma})=\Gamma$. 
Hence by Theorem \ref{indisoinn}, we can study the cohomology of a infra-solvmanifolds $M_{\Gamma}$.

For a  torsion-free virtually polycyclic group $\Gamma$ and a full representation $\rho:\Gamma\to G$,
take  a splitting $G=T\ltimes U_{\Gamma}$ for a maximal reductive subgroup $T$.
Denote by $\frak u$  the Lie algebra of  the unipotent radical $U$ of $G$.
Then by Theorem \ref{indisoinn}, the cohomology of the DGA 
$\left(\bigwedge {\frak u}^{\ast}\right)^{T}$ is isomorphic to $H^{\ast}(M_{\Gamma},\Q)$.
By this isomorphism, we can find an integral symplectic form on the compact infra-solvmanifold $M_{\Gamma}$.

\begin{proposition}
If there exists a $2$-form $\omega\in \left(\bigwedge^{2} {\frak u}^{\ast}\right)^{T}$ such that $d\omega=0$ and $\omega$ is non-degenerate,
 then the   compact infra-solvmanifold $M_{\Gamma}$ admits a integral symplectic form.
\end{proposition}
\begin{proof}
We suppose that ${\rm rank}\, \Gamma=2n$.
It is sufficient to show that $M_{\Gamma}$ admits a symplectic form $\alpha$ such that $[\alpha]\in H^{2}(M_{\Gamma},\Q)$.
It is known that if we have $a\in H^{2}(M_{\Gamma},\R)$ so that $a^{n}\not=0$, then we have a symplectic form $\alpha$ which is a representative of $a$ (see \cite{Ksy}).

Now we suppose that
 there exists a $2$-form $\omega\in \left(\bigwedge^{2} {\frak u}^{\ast}\right)^{T}$ such that $d\omega=0$ and $\omega$ is non-degenerate.
Consider the cohomology class $[\omega]\in H^{2}(M_{\Gamma},\Q)$.
Then, since $\omega\in \left(\bigwedge^{2} {\frak u}^{\ast}\right)^{T}$ is non-degenerate, we have $[\omega]^{n}\not=0$.
Hence, taking a symplectic form which is a representative of $[\omega]$, we can prove the proposition.
\end{proof}

\begin{example}
Let $A$ be a semi-simple matrix such that $A\in SL_{m}(\Z)$ and for any non-zero integer $n$ we suppose $A^{n}\not=1$ where $1$ is the unit matrix.
Consider the semi-direct product $\Gamma=\Z\ltimes \Z^{m}$ such that 
the action of $\Z$ on $\Z^{m}$ is given by 
$\Z\ni t\mapsto A^{t}\in {\rm Aut}(\Z^{m})$.
The group $\Gamma$ is torsion-free polycyclic of rank $m+1$.
Take $B=\left(
\begin{array}{ccccccc}
1&0\\
0&A
\end{array}
\right) \in  SL_{m+1}(\Z)$.
Let $T$ be the Zariski-closure of $\langle B\rangle $ in $SL_{m+1}(\Q)$.
By the assumption, $T$ is diagonalizable.
consider the $\Q$-algebraic group $G=T\ltimes \Q^{m+1}$ with the unipotent radical $\Q^{m+1}$.
We have the injective homomorphism
\[\rho: \Gamma \ni (t,v)\mapsto (B^{t},t,v)\in G
\]
where $t\in \Z$ and $v\in \Z^{m}$.
Then we can easily check that $\rho(\Gamma)$ is Zariski-dense in $G$ and so $\rho$ is full.  
Thus by  Theorem \ref{indisoinn}, we have an isomorphism
\[H^{\ast}(M_{\Gamma},\Q)\cong \left(\bigwedge \Q^{m+1}\right)^{T}=\left(\bigwedge \Q^{m+1}\right)^{\langle B\rangle}.
\]

Additionally, we assume that 
\[A=\left(
\begin{array}{ccccccc}
1&&&\\
&A_{1}&&\\
&&\ddots&\\
&&&A_{k}
\end{array}\right)
\]
such that $A_{1},\dots, A_{k}\in SL_{2}(\Z)$ and they are semi-simple.
By the above assumption, some $A_{i}$ satisfies $A^{n}_{i}\not=1$ for any non-zero integer $n$.
Then, taking the standard basis $e_{1}\dots e_{2k+2}$ of $\Q^{2k+2}$,
we obtain the non-degenerate two form $\omega\in  \left(\bigwedge^{2} \Q^{2k+2}\right)^{\langle B\rangle}$
such that 
\[\omega=e_{1}\wedge e_{2}+\dots +e_{2k+1}\wedge e_{2k+2}.
\]
Hence in this case $M_{\Gamma}$ admits a integral  symplectic form $\omega$.
By  \cite[Proposition 1.1, Proposition 1.4]{Kfo}, we can say that $M_{\Gamma}$ is formal in the sense of Sullivan (\cite{DGMS}, \cite{Sul}) and $(M_{\Gamma},\omega)$ satisfies the hard Lefschetz property.
On the other hand, since the matrix $A$ does not have a finite period, $\Gamma$ is not virtually nilpotent and hence by the result in \cite{AN} $M_{\Gamma}$ does not admit a K\"ahler structure.
Using integral symplectic form $\omega$, we construct the following examples:
\begin{newexample}

Since $M_{\Gamma}$ is $2k+2$-dimensional manifold with integral symplectic form $\omega$,
we have a symplectic emmbedding $M_{\Gamma}\to \C P^{N}$ for $N\ge 2k+3$ (\cite{Gro}, \cite{Ts}).
Hence, as McDuff's construction \cite{Mcd}, we obtain the symplectic blow-up $X_{\Gamma}$ of $\C P^{N}$ along $M_{\Gamma}$.
By \cite[Theorem 2.2]{Cav}, the symplectic blow-up $X_{\Gamma}$ satisfies the hard Lefschetz property.
Moreover, in \cite{LS}, it is announced that the blow-up along a manifold $M$ symplectically embedded in a large enough complex projective space is formal if and only if $M$ is formal.
If this fact true, then $X_{\Gamma}$ is formal and hence we obtain a formal and hard Lefschetz symplectic manifolds such that we do not know whether they admit K\"ahler structures.
\end{newexample}
\begin{newexample}
By the integral symplectic form $\omega$ on $M_{\Gamma}$, 
we obtain the principal circle bundle $P_{\Gamma}\to M_{\Gamma}$ associated with $[\omega]\in H^{2}(M_{\Gamma},\Z)$.
It is known that $P_{\Gamma}$ is a regular contact manifold.
Recently, the hard Lefschetz property on contact manifolds is defined. (see \cite{CNY}, \cite{Li})
In \cite{Li}, it is proved that the contact manifold $P_{\Gamma}$ satisfies hard Lefschetz property if and only if the symplectic manifold $M_{\Gamma}$ satisfies hard Lefschetz property.
Hence $P_{\Gamma}$ satisfies hard Lefschetz property.
In \cite{KSS}, it is proved that polycyclic fundamental groups of  compact Sasakian manifolds are virtually nilpotent.
Since $\Gamma$ is not virtually nilpotent, the fundamental group of $P_{\Gamma}$ is  not virtually nilpotent.
Hence $P_{\Gamma}$ does not admit a Sasakian structure.

\end{newexample}
\end{example}

\end{document}